OpenStudy (pphalke):

A company paid $225,000 in sales tax to the government from its auction sales last year, but later found out it over paid 25%. How much money should the company claim back from the government?

OpenStudy (danielle13asks):

Sorry I don't think I can help! But Love your support towards Christina Grimmie!

OpenStudy (hollowdensity):

@pphalke you need to find 25% of the said tax. '%' means or has the value of 0.01.

OpenStudy (pphalke):

@HollowDensity ok hw do i go about this process?

OpenStudy (hollowdensity):

u should think of it like this. let 'x' be the required value. then x corresponds to 225,000 just as 25 corresponds to 100.

OpenStudy (hollowdensity):

which mathematically translates to (x/225000) = (1/100)

OpenStudy (hollowdensity):

@pphalke did it make some sense to you?

OpenStudy (pphalke):

yes it made a lot of sense @HollowDensity

OpenStudy (hollowdensity):

@pphalke sorry, made mistake. (x/225000) = (25/100)

OpenStudy (hollowdensity):

its 25 not 1

OpenStudy (mathstudent55):

The calculation above will not give you the correct answer. $225,000 is 25% more than the correct amount of taxes the company owed. Taking 25% off $225,000 will not give you that amount.

OpenStudy (pphalke):

@mathstudent55 i think its supposed to be 45000

OpenStudy (hollowdensity):

@mathstudent55 my bad. i re-read the question and u r correct.

OpenStudy (mathstudent55):

The above calculation gives you $56,250 which is incorrect. Let's state this problem correctly. The company should have paid x amount of dollars in taxes. That means that x amount is 100% of the taxes. The company mistakenly paid 25% more than it should have, so the company paid 125% of the taxes is owed.

OpenStudy (mathstudent55):

125% of x is $225,000. What is x? 125% * x = 225,000 1.25x = 225,000 x = 180,000 The company should have paid $180,000 in taxes. Since it paid 25% extra, it paid 25% of $180,000 = 0.25 * $180,000 = $45,000

OpenStudy (hollowdensity):

@pphalke u can still apply the statement i mentioned above. suppose x is the amount that should have been paid then. x corresponds to 100 just as 225000 corresponds to 125.

OpenStudy (hollowdensity):

then difference of what is paid and what should have paid is the required value or amount..

OpenStudy (mathstudent55):

In percent problems, you need to be careful with taking the percent of the beginning amount or of the end mount. You must read the problem carefully to determine which percent to take.

OpenStudy (hollowdensity):

my head spinning. follow @mathstudent55. i messed it up again.

OpenStudy (mathstudent55):

No, you are correct now. "x is to 100 as 225,000 is 125" is correct. \(\dfrac{x}{100} = \dfrac{225000}{125} \) 125x = 22500000 x = 180000 Then you can do the subtraction you mentioned: 225000 - 180000 = 45000